OpenStudy (anonymous):

What is the inverse of y=6x-2

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow y +2=6x }\) \(\Large \color{Black}{\Rightarrow x = {y + 2 \over 6} }\) So we can say that the inverse is (x + 2)/6

OpenStudy (anonymous):

y = 6x - 2 x = 6y - 2 x + 2 = 6y (x + 2)/6 = y The inverse is: y = (x + 2)/6